# Text-to-Image Generation Via Energy-Based CLIP

### v1

@@ 3.2 Contrastive Adversarial Training
Similar to CLIPAG (Ganz and Elad 2024), we extend adversarial training to the contrastive loss, forming a contrastive adversarial loss. Given an input batch of image-text pairs, we perform a two-step procedure – (i) crafting visual adversarial examples that maximizes the similarity w.r.t the matching texts; (ii) optimizing the vision encoder’s weight to minimize the contrastive adversarial loss. The adversarial contrastive loss matrix is illustrated on the lower-left side of Figure 3. In this matrix, the vertical green vector represents the encodings of the adversarial visual inputs, while the horizontal pink vector represents the textual encodings. This approach ensures that the model learns to be robust against adversarial perturbations and equips it with semantically meaningful gradients (i.e., PAG), which enhance the iterative generation process.
@@ 3.3 Training Procedure
CLIP-JEM’s objective is the combination of contrastive adversarial and contrastive energy losses, and the overall training protocol is described in the supplementary materials (algorithm 1). The goal of CLIP-JEM is to obtain an image-text energy model capable of generating images based on textual descriptions. On the right side of Figure 3, we illustrate this generation process (piece-wise linear arrow), starting from a random sample from a canonical distribution (marked in gray). In this figure, any image-text pair is represented by two values: CLIP score and perceptual quality. The former measures the alignment of the pair according to our CLIP model, while the latter is a conceptual measure of the image’s visual quality. We aim to generate samples with high CLIP scores (low energy) and good perceptual quality. We analyze the contribution of both objectives to this goal. The contrastive energy loss prevents CLIP-JEM from assigning high CLIP scores to low-quality images, as throughout the energy training, the model is trained to lower the CLIP scores of “negative” samples and provide high scores to the “positive” ones, which have high perceptual quality. The contrastive adversarial objective trains the model against adversarial inputs, namely, imperceptible visual changes that significantly decrease the CLIP score (these samples reside in the bottom right corner of the figure’s center part). Thus, this loss prevents CLIP-JEM from assigning low CLIP scores to high-quality inputs. Overall, combining the two objectives eliminates the red modes in the upper-right panel of Figure 3, preventing the iterative generation from drawing such samples. This, in return, leads to drawing samples of the green mode, which are realistically looking images that align with their corresponding text. We demonstrate the contribution of combining these two objectives in the appendix D.
@@ 4 Experiments
We train different variants of CLIP using Algorithm 1 – including ViT-B/32 and ConvNext in base, large, and XXL configurations on the extensive image-caption DataComp dataset (Gadre et al. 2024) for $20,000$ steps. Throughout the training process, we keep the text encoder frozen and update solely the vision encoder. Implementation and training details are provided in the supplementary materials. To analyze the performance of CLIP-JEM, we first evaluate it in the text-to-image generation setting Section 4.1). Next, we demonstrate its effectiveness as a guiding model (Section 4.2). Lastly, we show that CLIP-JEM can serve as an improved evaluation metric compared to the “vanilla” CLIP, attributed to its robustness and awareness of perceptual quality.
@@ 4.1 Text-To-Image Generation
Similar to the training procedure, we perform pixel-space optimization to generate samples. Given a target prompt, we initialize the image as a random sample from a uniform distribution and perform $50$ steps to maximize the cosine similarity with respect to the text, using an AdamW optimizer with no momentum. We evaluate the performance of CLIP-JEM in two main setups: image quality and compositionality.
@@
We compare CLIP-JEM with other generative models using T2I CompBench (Huang et al. 2023), which evaluates open-world compositional text-to-image generation across attribute binding (color, shape and texture), object relationship (spatial and non-spatial) and complex. Despite advances in text-to-image generation, models still struggle to compose objects with different characteristics and relationships into a coherent image. Following CompBench procedure, we generate $10$ samples per prompt and average the results on the validation sets of each category, using the same evaluation metrics as in the original paper (B-VQA, UniDet, CLIP, and 3-in-1 for the attribute binding, spatial, non-spatial, and complex categories). Table 2 reports CLIP-JEM ’s results compared to top-performing models, including an average score across six categories. CLIP-JEM excels in attribute binding, associating attributes with corresponding objects in generated images, and performs well in the non-spatial relationship category (e.g., “speak to” and “look at”). However, it scores lower in the spatial relationship category due to CLIP’s known limitations in spatial compositionality (Yuksekgonul et al. 2022; Lewis et al. 2022). In the complex category, involving multiple objects and attributes, CLIP-JEM performs well despite containing spatial relationships, due to its strengths in attribute binding and non-spatial understanding. Overall, CLIP-JEM outperforms most baselines in compositional generation, despite being smaller. Specifically, it surpasses methods deliberately designed to tackle compositionality and rely on a much stronger generative model (Stable Diffusion v2). We attribute this success to the generative and discriminative objectives combination, enabling effective alignment with compositional prompts.
@@
Stable v1-4 0.3765 0.3576 0.4156 0.1246 0.3079 0.3080 0.3150
Stable v2 0.5065 0.4221 0.4922 0.1342 0.3127 0.3386 0.3677
Composable + SD v2 0.4063 0.3299 0.3644 0.0800 0.2980 0.2898 0.2947
Structured + SD v2 0.4990 0.4218 0.4900 0.1386 0.3111 0.3355 0.3660
Attn-Exct + SD v2 0.6400 0.4517 0.5963 0.1455 0.3109 0.3401 0.4141
@@ 4.2 Text Guidance Using CLIP-JEM
As shown in fig. 1, CLIP-JEM possesses semantically meaningful gradients with respect to a given text. In this section, we demonstrate the guidance capability of our method in two main settings: diffusion guidance and improving CLIP-based generative frameworks, utilizing a ConvNext Large model.
@@
We utilize CLIP-JEM as a guiding technique to transform unconditional diffusion models trained on ImageNet (Dhariwal and Nichol 2021; Ahn et al. 2024) into text-conditioned ones using only $25$ DDIM steps $25$ (Song, Meng, and Ermon 2020). In each DDIM step $t$, we update the estimations for the clean image $\hat{\mathbf{x}}_{0}$ using the gradients of CLIP-JEM (eq. 5) and use it to compute $\mathbf{x}_{t-1}$ and continue the reverse DDIM process.
@@
We provide qualitative results of text-based diffusion guidance in Figure 4, showcasing the capability of CLIP-JEM to convert unconditional diffusion models to text-based ones, thereby enhancing their utility.
@@
CLIP is widely used in text-to-image generation frameworks to update the resulting image to better align with the target text in CLIP’s space. This is typically done using the input gradients of the CLIP’s vision encoder with respect to the text (maximizing the cosine similarity). However, these gradients tend to be semantically meaningless, as demonstrated in Figure 1. To mitigate this, a multiview augmentation pipeline is often employed to acquire semantically meaningful gradients. Our method, on the other hand, inherently produces gradients that convey rich semantic information. Consequently, using CLIP-JEM can eliminate the need for multiview augmentations, thereby improving computational efficiency. To demonstrate this, we experiment with two such frameworks: CLIPDraw (Frans, Soros, and Witkowski 2022) and VQGAN+CLIP (Crowson et al. 2022). CLIPDraw generates drawings by optimizing the parameters of Bézier curves using CLIP, and VQGAN+CLIP updates the latent code of a VQGAN to enable text-to-image generation. In both cases, we do not perform multiview augmentations and report the results with augmentations in the supplementary materials. In Figure 5, we present the results of CLIPDraw on target texts from the original paper, along with CLIP’s top prediction for each prompt. As shown, using CLIP-JEM leads to significantly better visual results, which are more aligned with the text than those generated by the “vanilla” CLIP. The high percentages in the top predictions imply that the images generated by CLIP have an adversarial nature, maximizing the score without performing significant modifications. Next, we evaluate the effect of using our approach to guide VQGAN. To this end, we prompted ChatGPT to provide artistic target texts, aligning with the original paper’s domain. As seen in Figure 6, removing the augmentation pipeline results in non-meaningful outputs, whereas CLIP-JEM generates semantically meaningful images. These results strongly attest to the improved guidance capabilities of CLIP-JEM.
Figure 5: CLIPDraw results
Figure 6: VQGAN+CLIP results.
@@ 4.3 CLIP-JEM as an Evaluation Metric
Table 3: Robustness To Adversarial Perturbation.
CLIP is often used to evaluate text-to-image generative tasks by measuring cosine similarity between textual descriptions and images in its embedding space, known as CLIP-T. We compare CLIP-JEM with the standard CLIP model, focusing on the ViT-B/32 commonly used for this purpose. Using TEdBench (Kawar et al. 2023), we evaluate CLIP-T scores for various inputs, including outputs from a top-performing generative model (Imagic), source images (“No Edit”), and “Noise” images. Results are shown in table 3 under “Vanilla”. We also assess robustness under adversarial attacks with a low perturbation budget ($\epsilon=\nicefrac{{2}}{{255}}$). These attacks aim to increase scores for “bad” images and decrease scores for “good” ones. Our findings indicate that CLIP is highly susceptible to adversarial attacks, resulting in higher CLIP-T scores for non-edited and noise images than for Imagic outputs. In contrast, CLIP-JEM remains robust, maintaining higher scores for Imagic outputs even under adversarial attacks.
We further analyze the effect of perceptual quality on CLIP-T scores by blending Imagic images ($\mathbf{x}_{Imagic}$) with uniform noise: $\lambda\mathbf{x}_{Imagic}+(1-\lambda)\mathbf{u}$ where $\mathbf{u}\sim\mathcal{U}[0,1].$ We plot CLIP-T scores for varying $\lambda$ values in Figure 7, normalizing scores to $1.0$ at $\lambda=0$. While the standard CLIP model prefers noisier versions up to $\lambda=0.45$, CLIP-JEM ’s scores decrease with increasing noise, indicating greater sensitivity to image quality, attributed to the contrastive energy objective making the model to assign high energy to non-real images.
@@ Appendix A Detailed background on energy-based models
Energy-based models (EBMs) (LeCun et al. 2006) are a class of probabilistic models that define a probability distribution over the data points via an energy function. Specifically, EBMs utilize the fact that any probability density $p(\mathbf{x})$ for $\mathbf{x}\in\mathbb{R}^{D}$ can be written as
@@
where $E_{\theta}(\cdot)$ is the energy function, which assigns scalar values to data points ($E_{\theta}:\mathbb{R}^{D}\rightarrow\mathbb{R}$) and $Z(\theta)$ is a normalizing constant. In EBMs, the probability of a given data point is determined by its energy, where lower energy indicates a higher probability, as can be seen in Equation 6.
@@ Appendix B Implementation details
We implement our method upon the OpenClip codebase [^3] https://github.com/mlfoundations/open˙clip. We consider the following model architectures from the model zoo – convnext_xxlarge, convnext_large_d, convnext_base_w and ViT-B-32 with the following pretrained weights, respectively – laion2b_s34b_b82k_augreg_soup, laion2b_s26b_b102k_augreg, laion2b_s13b_b82k_augreg and openai. In table 4, we report the training hyperparameters for CLIP-JEM . We use these hyperparameters to train the different architectures on DataComp. During training, in the generation process of the “negative” samples we employ a momentum of $0.9$ to the AdamW optimizer. However, throughout the inference phase, we do not utilize momentum at all. We aim to make our code and pretrained models publicly available upon acceptance.
@@
To measure the quality and fidelity of the generated images of our method, we compare it to strong baselines using MS-COCO dataset (table 1). Specifically, we compare CLIP-JEM to text-to-image generative models of different types: (i) GAN-based – Stack-GAN (Zhang et al. 2017), AttnGAN (Xu et al. 2017), LAFITE (Zhou et al. 2022), StyleGAN-T (Sauer et al. 2023), and GigaGAN (Kang et al. 2023b) (ii) Diffusion-based – GLIDE (Nichol et al. 2022), and LDM (Rombach et al. 2022c), and Autoregressive ones – DALL-E (Ramesh et al. 2021), CogView (Ding et al. 2021), and NÜWA (Wu et al. 2021). Our reported results and model sizes originate from the respective papers. The CLIPAG baseline results were obtained by us, by removing the contrastive energy loss term, using the same architecture and hyperparameters. As for the CompBench results (table 2), we report the one from the benchmark’s paper.
As for the improving CLIP-based generative frameworks experiments, we use an ADM-based model (Dhariwal and Nichol 2021) with perturbed-attention guidance (Ahn et al. 2024) which strongly improves the unconditional generation as our baseline for diffusion guidance.
@@
We train 4 different architectures of CLIP-JEM. As different model size and structure affects the runtime complexity, we report the time of our generation process using a batch size of $1$ using an Nvidia A40 GPU in table 5. As expected, increasing the model size leads to more memory consumption and increases the time per sample. However, the ConvNext-B generation time is slightly larger than the ConvNext-L. This is due to the fact the we utilize the wide variant of the ConvNext-B, which is not available in ConvNext-L.

### v2

@@ 3.2 Contrastive Adversarial Training
Similar to CLIPAG (Ganz & Elad, 2024), we extend adversarial training to the contrastive loss, forming a contrastive adversarial loss. Given an input batch of image-text pairs, we perform a two-step procedure – (i) crafting visual adversarial examples that maximizes the similarity w.r.t the matching texts; (ii) optimizing the vision encoder’s weight to minimize the contrastive adversarial loss. The adversarial contrastive loss matrix is illustrated on the lower-left side of Figure˜3, and in eq.˜5. In this matrix, the vertical green vector represents the encodings of adversarial visual input, while the horizontal pink vector represents the textual encodings. This approach ensures that the model learns to be robust against adversarial perturbations and equips it with semantically meaningful gradients (i.e., PAG), which enhance the iterative generation process.
$\min_{\theta_{I},\theta_{T}}\sum_{(\mathbf{x},t)\in\mathcal{D}}\max_{\delta\in\Delta}(1-\mathcal{\operatorname{CosineSimilarity}}(f^{I}_{\theta_{I}}(\mathbf{x}+\delta),f^{T}_{\theta_{T}}(t))).$ (5)
@@ 3.3 Training Procedure
CLIP-JEM’s objective is the combination of contrastive adversarial and contrastive energy losses, and the overall training protocol is described in the supplementary materials (algorithm˜1). The goal of CLIP-JEM is to obtain an image-text energy model capable of generating images based on textual descriptions. On the right side of Figure˜3, we illustrate this generation process (piece-wise linear arrow), starting from a random sample from a canonical distribution (marked in gray). In this figure, any image-text pair is represented by two values: CLIP score and perceptual quality. The former measures the alignment of the pair according to our CLIP model, while the latter is a conceptual measure of the image’s visual quality. We aim to generate samples with high CLIP scores (low energy) and good perceptual quality. We analyze the contribution of both objectives to this goal. The contrastive energy loss prevents CLIP-JEM from assigning high CLIP scores to low-quality images, as throughout the energy training, the model is trained to lower the CLIP scores of “negative” samples and provide high scores to the “positive” ones, which have high perceptual quality. The contrastive adversarial objective trains the model against adversarial inputs, namely, imperceptible visual changes that significantly decrease the CLIP score (these samples reside in the bottom right corner of the figure’s center part). Thus, this loss prevents CLIP-JEM from assigning low CLIP scores to high-quality inputs. Overall, combining the two objectives eliminates the red modes in the upper-right panel of Figure˜3, preventing the iterative generation from drawing such samples. This, in return, leads to drawing samples of the green mode, which are realistically looking images that align with their corresponding text. We demonstrate the contribution of combining these two objectives in the appendix˜D.
@@ 4 Experiments
We train different variants of CLIP using Algorithm˜1 – including ViT-B/32 and ConvNext in base, large, and XXL configurations on the extensive image-caption DataComp dataset (Gadre et al., 2024) for 20,00020,00020 , 000 steps. Throughout the training process, we keep the text encoder frozen and update solely the vision encoder. Implementation and training details are provided in the supplementary materials. To analyze the performance of CLIP-JEM, we first evaluate it in the text-to-image generation setting Section˜4.1). Next, we demonstrate its effectiveness as a guiding model (Section˜4.2). Lastly, we show that CLIP-JEM can serve as an improved evaluation metric compared to the “vanilla” CLIP, attributed to its robustness and awareness of perceptual quality.
@@ 4.1 Text-To-Image Generation
Similar to the training procedure, we perform pixel-space optimization to generate samples. Given a target prompt, we initialize the image as a random sample from a uniform distribution and perform 505050 steps to maximize the cosine similarity with respect to the text, using an AdamW optimizer with no momentum. We evaluate the performance of CLIP-JEM in two main setups: image quality and compositionality.
@@
We compare CLIP-JEM with other generative models using T2I CompBench (Huang et al., 2023), which evaluates open-world compositional text-to-image generation across attribute binding (color, shape and texture), object relationship (spatial and non-spatial) and complex. Despite advances in text-to-image generation, models still struggle to compose objects with different characteristics and relationships into a coherent image. Following CompBench procedure, we generate 101010 samples per prompt and average the results on the validation sets of each category, using the same evaluation metrics as in the original paper (B-VQA, UniDet, CLIP, and 3-in-1 for the attribute binding, spatial, non-spatial, and complex categories). Table˜2 reports CLIP-JEM ’s results compared to top-performing models, including an average score across six categories. CLIP-JEM excels in attribute binding, associating attributes with corresponding objects in generated images, and performs well in the non-spatial relationship category (e.g., “speak to” and “look at”). However, it scores lower in the spatial relationship category due to CLIP’s known limitations in spatial compositionality (Yuksekgonul et al., 2022; Lewis et al., 2022). In the complex category, involving multiple objects and attributes, CLIP-JEM performs well despite containing spatial relationships, due to its strengths in attribute binding and non-spatial understanding. Overall, CLIP-JEM outperforms most baselines in compositional generation, despite being smaller. Specifically, it surpasses methods deliberately designed to tackle compositionality and rely on a much stronger generative model (Stable Diffusion v2). We attribute this success to the generative and discriminative objectives combination, enabling effective alignment with compositional prompts.
@@
SD1.4 0.3765 0.3576 0.4156 0.1246 0.3079 0.3080 0.3150
SD2 0.5065 0.4221 0.4922 0.1342 0.3127 0.3386 0.3677
Composable (SD2) 0.4063 0.3299 0.3644 0.0800 0.2980 0.2898 0.2947
Structured (SD2) 0.4990 0.4218 0.4900 0.1386 0.3111 0.3355 0.3660
Attn-Exct (SD2) 0.6400 0.4517 0.5963 0.1455 0.3109 0.3401 0.4141
CLIP-JEM
@@ 4.2 Text Guidance Using CLIP-JEM
As shown in fig.˜1, CLIP-JEM possesses semantically meaningful gradients with respect to a given text. In this section, we demonstrate the guidance capability of our method in two main settings: diffusion guidance and improving CLIP-based generative frameworks, utilizing a ConvNext Large model.
@@
We utilize CLIP-JEM as a guiding technique to transform unconditional diffusion models trained on ImageNet (Dhariwal & Nichol, 2021; Ahn et al., 2024) into text-conditioned ones using only 252525 DDIM steps 252525 (Song et al., 2020). In each DDIM step $t$, we update the estimations for the clean image $\hat{\mathbf{x}}_{0}$ using the gradients of CLIP-JEM (eq.˜6) and use it to compute $\mathbf{x}_{t-1}$ and continue the reverse DDIM process.
@@
We provide qualitative results of text-based diffusion guidance in Figure˜4, showcasing the capability of CLIP-JEM to convert unconditional diffusion models to text-based ones, thereby enhancing their utility.
@@
CLIP is widely used in text-to-image generation frameworks to update the resulting image to better align with the target text in CLIP’s space. This is typically done using the input gradients of the CLIP’s vision encoder with respect to the text (maximizing the cosine similarity). However, these gradients tend to be semantically meaningless, as demonstrated in Figure˜1. To mitigate this, a multiview augmentation pipeline is often employed to acquire semantically meaningful gradients. Our method, on the other hand, inherently produces gradients that convey rich semantic information. Consequently, using CLIP-JEM can eliminate the need for multiview augmentations, thereby improving computational efficiency. To demonstrate this, we experiment with two such frameworks: CLIPDraw (Frans et al., 2022) and VQGAN+CLIP (Crowson et al., 2022). CLIPDraw generates drawings by optimizing the parameters of Bézier curves using CLIP, and VQGAN+CLIP updates the latent code of a VQGAN to enable text-to-image generation. In both cases, we do not perform multiview augmentations and report the results with augmentations in the supplementary materials. In Figure˜5(a), we present the results of CLIPDraw on target texts from the original paper, along with CLIP’s top prediction for each prompt. As shown, using CLIP-JEM leads to significantly better visual results, which are more aligned with the text than those generated by the “vanilla” CLIP. The high percentages in the top predictions imply that the images generated by CLIP have an adversarial nature, maximizing the score without performing significant modifications. Next, we evaluate the effect of using our approach to guide VQGAN. To this end, we prompted ChatGPT to provide artistic target texts, aligning with the original paper’s domain. As seen in Figure˜5(b), removing the augmentation pipeline results in non-meaningful outputs, whereas CLIP-JEM generates semantically meaningful images. These results strongly attest to the improved guidance capabilities of CLIP-JEM.
(a) CLIPDraw results
(b) VQGAN+CLIP results
Figure 5: Improving CLIP-Based Generative Frameworks via CLIP-JEM. Qualitative results of CLIP-JEM compared to CLIP using CLIPDraw and VQAGAN+CLIP in a zero-shot setting.
@@ 4.3 CLIP-JEM as an Evaluation Metric
CLIP is often used to evaluate text-to-image generative tasks by measuring cosine similarity between textual descriptions and images in its embedding space, known as CLIP-T. We compare CLIP-JEM with the standard CLIP model, focusing on the ViT-B/32 commonly used for this purpose. Using TEdBench (Kawar et al., 2023), we evaluate CLIP-T scores for various inputs, including outputs from a top-performing generative model (Imagic), source images (“No Edit”), and “Noise” images. Results are shown in section˜4.3 under “Vanilla”. We also assess robustness under adversarial attacks with a low perturbation budget ($\epsilon=\nicefrac{{2}}{{255}}$). These attacks aim to increase scores for “bad” images and decrease scores for “good” ones. Our findings indicate that CLIP is highly susceptible to adversarial attacks, resulting in higher CLIP-T scores for non-edited and noise images than for Imagic outputs. In contrast, CLIP-JEM remains robust, maintaining higher scores for Imagic outputs even under adversarial attacks.
We further analyze the effect of perceptual quality on CLIP-T scores by blending Imagic images ($\mathbf{x}_{Imagic}$) with uniform noise: $\lambda\mathbf{x}_{Imagic}+(1-\lambda)\mathbf{u}$ where $\mathbf{u}\sim\mathcal{U}[0,1].$ We plot CLIP-T scores for varying $\lambda$ values in Figure˜6, normalizing scores to 1.01.01.0 at $\lambda=0$. While the standard CLIP model prefers noisier versions up to $\lambda=0.45$, CLIP-JEM ’s scores decrease with increasing noise, indicating greater sensitivity to image quality, attributed to the contrastive energy objective making the model to assign high energy to non-real images.
Table 3: Robustness To Adversarial Perturbation.
@@ Appendix A Detailed background on energy-based models
Energy-based models (EBMs) (LeCun et al., 2006) are a class of probabilistic models that define a probability distribution over the data points via an energy function. Specifically, EBMs utilize the fact that any probability density $p(\mathbf{x})$ for $\mathbf{x}\in\mathbb{R}^{D}$ can be written as
@@
where $E_{\theta}(\cdot)$ is the energy function, which assigns scalar values to data points ($E_{\theta}:\mathbb{R}^{D}\rightarrow\mathbb{R}$) and $Z(\theta)$ is a normalizing constant. In EBMs, the probability of a given data point is determined by its energy, where lower energy indicates a higher probability, as can be seen in Equation˜7.
@@ Appendix B Implementation details
We implement our method upon the OpenClip codebase [^3] https://github.com/mlfoundations/open_clip. We consider the following model architectures from the model zoo – convnext_xxlarge, convnext_large_d, convnext_base_w and ViT-B-32 with the following pretrained weights, respectively – laion2b_s34b_b82k_augreg_soup, laion2b_s26b_b102k_augreg, laion2b_s13b_b82k_augreg and openai. In table˜4, we report the training hyperparameters for CLIP-JEM . We use these hyperparameters to train the different architectures on DataComp. To train our models, we use 888 A40 GPUs for training. Training the largest variant (ConvNext-XXL) for  [^20] K iterations takes 10 days. During training, in the generation process of the “negative” samples we employ a momentum of 0.90.90.9 to the AdamW optimizer. However, throughout the inference phase, we do not utilize momentum at all. We aim to make our code and pretrained models publicly available upon acceptance.
@@
To measure the quality and fidelity of the generated images of our method, we compare it to strong baselines using MS-COCO dataset (table˜1). Specifically, we compare CLIP-JEM to text-to-image generative models of different types: (i) GAN-based – Stack-GAN (Zhang et al., 2017), AttnGAN (Xu et al., 2017), LAFITE (Zhou et al., 2022), StyleGAN-T (Sauer et al., 2023), and GigaGAN (Kang et al., 2023b) (ii) Diffusion-based – GLIDE (Nichol et al., 2022), and LDM (Rombach et al., 2022c), and Autoregressive ones – DALL-E (Ramesh et al., 2021), CogView (Ding et al., 2021), and NÜWA (Wu et al., 2021). Our reported results and model sizes originate from the respective papers. The CLIPAG baseline results were obtained by us, by removing the contrastive energy loss term, using the same architecture and hyperparameters. As for the CompBench results (table˜2), we report the one from the benchmark’s paper.
As for the improving CLIP-based generative frameworks experiments, we use an ADM-based model (Dhariwal & Nichol, 2021) with perturbed-attention guidance (Ahn et al., 2024) which strongly improves the unconditional generation as our baseline for diffusion guidance.
@@
We train 4 different architectures of CLIP-JEM. As different model size and structure affects the runtime complexity, we report the time of our generation process using a batch size of 111 using an Nvidia A40 GPU in table˜5. As expected, increasing the model size leads to more memory consumption and increases the time per sample. However, the ConvNext-B generation time is slightly larger than the ConvNext-L. This is due to the fact the we utilize the wide variant of the ConvNext-B, which is not available in ConvNext-L.